OpenStudy (anonymous):

whats the improper fraction of 3 1/6?

OpenStudy (anonymous):

19/6

OpenStudy (anonymous):

u sure

OpenStudy (anonymous):

Why wouldn't I be sure?

OpenStudy (anonymous):

It is 19/6

OpenStudy (asnaseer):

think of it like this:\[3\frac{1}{6}=3+\frac{1}{6}=\frac{18}{6}+\frac{1}{6}=\frac{19}{6}\]